An electron is continuously accelerated in a vacuum tube  by applying a potential difference. If the de Broglie wavelength is decreased by 10%, the change in kinetic energy of the electron is
 

Grade:

1 Answers

Sumit Majumdar IIT Delhi
9 years ago
Dear student,
We have the equation being given by:
\lambda =\frac{h}{\sqrt{2mK.E}}=\frac{h}{\sqrt{2m}}\left ( K.E \right )^{-\frac{1}{2}}
So, we get:

\left (\frac{d\lambda}{\lambda} \right ) =\frac{-1}{2}}\left (\frac{dK.E.}{K.E} \right )
So,
\left (\frac{dK.E.}{K.E.} \right )\times 100 =2\left (\frac{d\lambda.}{\lambda} \right )\times 100=20%
Regards
Sumit